### v1

# line of sight CSI channel state information

[…]
LoS
line of sight
CSI
channel state information
NLoS
non line of sight
InH
indoor hotspot
InF
indoor factory
UMi
urban microcell
SHF
super high frequency
DS
delay spread
RMS
root mean square
AS
angular spread
DSS
dynamic spectrum sharing
3GPP
$3^{\text{rd}}$ Generation Partnership Program
NGJ-MB
next generation jammer mid-band
RF
radio frequency
ACLR
adjacent channel leakage power ratio
FCC
Federal Communications Commission
NTIA
National Telecommunications and Information Administration
CFO
carrier frequency offset
CRB
Cramér-Rao bound
AWGN
additive white Gaussian noise
FFT
fast Fourier transform
CP
cyclic prefix
MSE
mean squared error
MRC
maximum ratio combining
CPE
customer-premises equipment
TRP
transmission and reception point
PSD
power spectral density
SE
spectral efficiency
AR
augmented reality
XR
extended reality
ISAC
integrated sensing and communications
BS
base station
FR1
frequency range 1
FR2
frequency range 2
FR3
frequency range 3
PLE
pathloss exponent
mmWave
milli-meter wave
gNB
gNodeB
NR
new radio
UE
user equipment
MC
mutual coupling
SINR
signal-to-interference-plus-noise ratio
ITU
International Telecommunication Union
MIMO
multiple-input and multiple-output
ADC
analog-to-digital converter
DAC
digital-to-analog converter
PA
power amplifier
O-RU
open radio unit
DoF
degrees-of-freedom
UL
uplink
DL
downlink
SHO
soft handover
CDMA
code-division multiple access
VR
virtual reality
QAM
quadrature amplitude modulation
LTE
long-Term Evolution
SNR
signal-to-noise ratio
RMT
random matrix theory
ELAA
extremely large aperture array
NF
near-field
FF
far-field
RT
ray tracing
AI
artificial intelligence
RT
real-time
RAN
radio access network
O-RAN
open radio access network
RIC
RAN intelligent controller
EC
ergodic capacity
AoA
angle-of-arrival
SAR
synthetic aperture radar
RIS
reconfigurable intelligent surface
ILFD
injection-locked frequency divider
TSPC
true single-phase clock
ET
extended target
UWB
ultra-wide band
S&C
sensing and communication
SB
spectrum block
mMIMO
massive MIMO
NB
narrowband
EIRP
equivalent isotropic radiated power
TDoc
technical document
RAN
radio access network
WG1
working group
CIF
close-in free space model with a frequency-weighted path loss exponent
TSG
technical specification group
ASA
angular spread of arrival
Upper Mid-Band Spectrum for 6G: Vision, Opportunity and Challenges
Ahmad Bazzi, Roberto Bomfin, Marco Mezzavilla, Sundeep Rangan, Theodore Rappaport, Marwa Chafii Ahmad Bazzi, Roberto Bomfin, and Marwa Chafii are with the Engineering Division, New York University Abu Dhabi (NYUAD), 129188, UAE (email: {ahmad.bazzi,roberto.bomfin, marwa.chafii}@nyu.edu)Ahmad Bazzi, Marwa Chafii, Sundeep Rangan, and Theodore Rappaport are with NYU WIRELESS, NYU Tandon School of Engineering, Brooklyn, 11201, NY, USA (email: {srangan,tsr}@nyu.edu).Marco Mezzavilla is with the Telecommunications Engineering at the Department of Electronics, Information, and Bioengineering (DEIB) at Politecnico di Milano, Milan, Italy. (email: marco.mezzavilla@polimi.it).
Abstract
Driven by the pursuit of gigabit-per-second data speeds for future 6G mobile networks, in addition to the support of sensing and artificial intelligence applications, the industry is expanding beyond crowded sub-6 GHz bands with innovative new spectrum allocations. In this paper, we chart a compelling vision for 6G within the frequency range 3 (FR3) spectrum, i.e. $7.125$-$24.25$ $\operatorname{GHz}$, by delving into its key enablers and addressing the multifaceted challenges that lie ahead for these new frequency bands. Here we highlight the physical properties of this never-before used spectrum by reviewing recent channel measurements for outdoor and indoor environments, including path loss, delay and angular spreads, and material penetration loss, all which offer insights that underpin future 5G/6G wireless communication designs. Building on the fundamental knowledge of the channel properties, we explore FR3 spectrum agility strategies that balance coverage and capacity (e.g. data rate) tradeoffs, while also examining coexistence with incumbent systems, such as satellites, radio astronomy, and earth exploration. Moreover, we discuss the potential of massive multiple-input multiple-output, compact and digital architectures, and evaluate the potential of multiband sensing for FR3 integrated sensing and communications. Finally, we outline 6G standardization features that are likely to emerge from 3GPP radio frame innovations and open radio access network developments.
Index Terms:
$6$G, upper midband, FR3, channel characteristics, spectrum agility, ISAC
I Introduction
By $2034$, global mobile data traffic is expected to grow by $5$x–$9$x, with artificial intelligence (AI) accounting for one-third of the traffic. There is an industry-wide consensus that $6$G is expected to launch around $2030$ with new spectrum in frequency range 3 (FR3) [1], which spans $7.125$-$24.25$ $\operatorname{GHz}$ [2]. To meet the $6$G timeline, it is crucial to allocate the necessary spectrum a few years in advance, in order to ensure incumbents are accommodated and that the technology is ready for mass adoption at the global launch of $6$G. Companies are actively conducting proof-of-concept product trials to address bottlenecks and tackle myriad implementation challenges. At the same time, regulators must determine how the spectrum will be allocated, and operators need to develop viable business cases and deployment/integration strategies for nationwide $6$G networks.
I-A The FR3 Spectrum
As the global wireless industry advances towards $6$G, the upper mid-band frequencies, being a part of the high super high frequency (SHF) bands, are critical resources due to their balance between coverage and bandwidth availability, compared to lower frequency range 1 (FR1) (up to $7.125\operatorname{GHz}$ [3]) and higher frequency range 2 (FR2) ranges, which includes $24.25\operatorname{GHz}$ to $71\operatorname{GHz}$ [4]. Often termed the “Golden Band”, or ”Goldilocks Spectrum”, for $6$G, FR3 frequencies ($7.125$-$24.25$ $\operatorname{GHz}$ [2] as depicted in Fig. 1 [1]) are particularly suited for enhancing network capacity while maintaining reasonable propagation characteristics, offering moderate propagation losses, enabling extensive urban and suburban reach using existing towers.
The FR3 spectrum holds growing importance and interest for industry and regulatory bodies like the U.S. National Telecommunications and Information Administration (NTIA), the $3^{\text{rd}}$ Generation Partnership Program (3GPP) [5] and the Federal Communications Commission (FCC) , who are currently evaluating its potential alongside existing mobile radio bands in order to expand cellular services. Besides mobile operator use, the upper mid-band is being considered to coexist with incumbent services such as satellite communications, radio astronomy, and earth exploration, in addition to warfare activities, such as the next generation jammer mid-band (NGJ-MB) operating within $509\operatorname{MHz}$ to $18\operatorname{GHz}$ band [6].
Figure 1: The spectrum showing the placement of FR3 Golden band, relative to its other counterparts according to different bodies and in various regions [1].
Notably, the NTIA highlighted specific FR3 bands, including $7.125\operatorname{GHz}$-$8.4\operatorname{GHz}$, and $14.8$-$15.35\operatorname{GHz}$, as part of the International Telecommunication Union (ITU) World Radio Conference $2023$ (WRC-$23$) agenda, emphasizing the pending authorization in the USA and hence the need for precise channel modeling to ensure effective spectrum utilization and sharing.
II FR3 Channel Characteristics
TABLE I: Omni PLEs from the CI PL model with 1 m reference distance, Omni DS, and Omni ASA at RX [1, 7, 8].
Env Freq. (GHz) Dist (m) LOS NLOS LOS Omni RMS DS NLOS Omni RMS DS LOS Omni RMS ASA NLOS Omni RMS ASA
Omni PLE Sigma Omni PLE Sigma
InH [1] 6.75 11–97 1.34 3.51 2.72 9.21 37.7 48 40.9 58.2
16.95 11–97 1.32 2.66 3.05 8.11 22.1 40.7 34.2 43.5
28 4–46 1.2 1.8 2.7 9.7 10.8 17.1 Med: 39.1 Med: 31.8
UMi [7] 6.75 40–1000 1.79 2.57 2.56 6.53 62.8 75.6 16.79 25.62
16.95 40–1000 1.85 4.05 2.59 8.78 46.5 65.8 18.23 19.08
28 31–187 2.02 8.98 3.56 8.91 26.7 46.3 Med:14 Med:30
InF [8] 6.75 9–38 1.39 1.86 1.78 2.46 14 30.3 23.71 66.38
16.95 9–38 1.75 3.1 2.11 3.29 12.7 29 14.32 50.4
II-A FR3 Channel Modeling
The channel model introduced in TR $38.901$ as part of the 3GPP global standard body designed to encompass the entire frequency range from $0.5$ to $100$ $\operatorname{GHz}$ [9]. Integrated channel models are indispensable for addressing diverse propagation scenarios, ranging from urban environments to free space, yet 3GPP developed the TR 38.901 channel models without many field measurements across much of the $100\operatorname{GHz}$ wide swath of the spectrum, and without any measurements from the FR3 bands. As a result, the model is an estimate, formulated from sparse measurements across particular bands within the $0.5$ to $100\operatorname{GHz}$ range. Channel modeling research for spectra above $6\operatorname{GHz}$ was completed in Release 17 of TR $38.901$ in April $2022$, offering a comprehensive set of models for evaluating various physical layer technologies. However, the primary focus of $5$G channel modeling has been on frequencies below $6\operatorname{GHz}$ (FR1) and above $24\operatorname{GHz}$ (FR3). To address this gap, frequency interpolation techniques were used by 3GPP to estimate channel parameters in the $7$-$24\operatorname{GHz}$ band. To establish a more accurate channel model for the FR3 band, it is crucial to validate the TR 38.901 model and carefully examine the details of FR3 channel parameterization. For this, true empirical characterization of the many channel parameters is required. 3GPP channel parameters include pathloss exponents, penetration losses, shadowing effects, cross-polarization discrimination (also referred to as cross-polarization isolation), in addition to delay and angular spreads across different environments, such as indoor hotspot (InH), indoor factory (InF) and urban microcell (UMi) under various conditions, such as line of sight (LoS) and non line of sight (NLoS) with minimal outages.
The findings of FR3 channel measurements conducted by NYU WIRELESS in [1, 7, 8] reveal that for wideband channels at $16.95\operatorname{GHz}$ in the FR3 band, the omnidirectional pathloss exponent (PLE) values, synthesized from directional channel measurements, in LoS scenarios were slightly lower than those in milli-meter wave (mmWave), indicating less signal attenuation and more of a waveguide effect over distance, e.g. the omnidirectional LoS PLEs in InH, UMi and InF for $6.75\operatorname{GHz}$, $16.75\operatorname{GHz}$, and $28\operatorname{GHz}$ are given in Table I, e.g. $1.34$ at $6.75\operatorname{GHz}$, $1.32$ at $16.75\operatorname{GHz}$, and $1.2$ at $28\operatorname{GHz}$ [1]. In UMi, the omnidirectional LoS PLEs are $1.79$ at $7\operatorname{GHz}$ [1], $1.85$ at $16.75\operatorname{GHz}$ [1], and $2.02$ at $28\operatorname{GHz}$ [1], which indicates slightly less loss over distance than free space channel (eg when PLE $=2$) and with less loss at lower frequencies. Although antenna patterns and gains vary widely, and higher frequencies enable the use of directional antennas that can offset channel loss in mmWave bands, 3GPP and researchers often use an omnidirectional channel model to standardize link analysis and ensure consistency in evaluations while enabling the use of any type of directional pattern to be applied to the models [10], [11]. NLoS PLEs were found to be higher than LoS scenario as is found in all 3GPP bands as per Table I, e.g. NLoS PLE for UMi of $2.56$ at $6.75\operatorname{GHz}$ [1], $2.59$ at $16.75\operatorname{GHz}$ [1], and for InF the NLoS PLE is $1.78$ at $6.75\operatorname{GHz}$ [8], and $2.11$ at $16.75\operatorname{GHz}$ [8], the loss over distance is much lower (eg $14.5\operatorname{dB}$ per decade less) than those observed at mmWave frequencies, e.g. NLoS PLE of $3.56$ at $28\operatorname{GHz}$, highlighting how FR3 offers improved coverage for a comparable equivalent isotropic radiated power (EIRP) and radio frequency (RF) bandwidth relative to higher frequencies. Referring to Table I, the PLE for LoS in UMi is $1.79$ for $6\operatorname{GHz}$ and $2.02$ for $28\operatorname{GHz}$, implying greater coverage with a $2.2\operatorname{dB}$ stronger signal per decade of distance, resulting in, $4.4\operatorname{dB}$ over $100\operatorname{m}$ and $6.6\operatorname{dB}$ stronger signal over $1\operatorname{km}$, which clearly demonstrates how a lower PLE translates into significant coverage improvement, which may be traded for higher capacity for users within a cell, a factor that will be of paramount importance to carriers when deploying $6$G technologies.
The time delay spread of a propagation channel has impact on signaling formats, in addition to accuracy for integrated sensing and communications (ISAC) applications such as position location and multi-user synchronization methods. Measurements in [1], [7], [8] show that root mean square (RMS) delay spread (DS) decreases with carrier frequency, with FR3 exhibiting smaller DS values than sub-$6$ $\operatorname{GHz}$, but larger than at FR2 mmWave frequencies, e.g. $14\operatorname{ns}$ and $12.7\operatorname{ns}$ LoS omnidirectional RMS DS at $6.75\operatorname{GHz}$ and $16.95\operatorname{GHz}$, respectively for InF. The range of RMS DS was found to be $12.7-37.7\operatorname{ns}$ across InF, UMi and InH. DS values observed in the field indicates that multipath components are closer in time at FR3, suggesting limited temporal dispersion which could favor high-speed, low-latency communications and more accurate position location and timing methods in dense environments. Furthermore, the measurements in [1] found that higher FR3 frequencies, i.e. $16.75\operatorname{GHz}$, exhibited a narrower RMS angular spread compared to the higher FR1 frequency at $6.75\operatorname{GHz}$, hence indicating fewer, more focused multipath components, which is beneficial for spatial multiplexing as it reduces interference between signal paths and allows more precise beamforming. Regarding material penetration, losses were consistently higher at $16.95\operatorname{GHz}$ in FR3 than at lower frequencies, with losses dependent on material type and polarization configuration, confirming the inherent limitations of FR3 in penetrating certain materials, such as low-emissivity (IRR) glass and concrete [1], yet able to allow more penetration that at FR3 mmWave [1].
When considering the design and comparison of air interface standards, such as the 3GPP TR 38.901 statistical channel model, the channel measurements [7, 8] conducted at new FR3 frequencies suggest that no major changes are needed in terms of numerology compared to existing 5G- new radio (NR) protocols; however, the beam management may need modifications due to the different temporal and spatial statistics and the different number of antenna elements at both gNodeB (gNB) and user equipment (UE). In particular, directional links rely on precise angular alignment of transmitter and receiver beams, a process facilitated through beam management operations. For next-generation cellular networks, mechanisms are being discussed in 3GPP technical documents that address beam management, such as R1111-$2409673$ in November $18-22$ as part of 3GPP technical specification group (TSG) radio access network (RAN) working group (WG1), in addition to other 3GPP contributions such as R1111-$2409710$, R1111-$2409741$ and more. Directional combining in beam management dramatically enhance link gains and coverage distances, as first shown in [12], thereby maintaining acceptable communication quality when beam combining is available – something that both FR2 and FR3 will support. However, as shown in [7] , as the carrier frequency increases, the angular spread tends to decrease. In addition, the empirical RMS angular spread of arrival (ASA) at both $6.75\operatorname{GHz}$ and $16.95\operatorname{GHz}$ was noticeably lower than the 3GPP model predictions [7]. Factors such as mobility, device rotation, and wide angular spread channels can impact beam alignment. In dynamic scenarios, beam measurements can quickly become outdated, leading to inaccurate beam selection and performance degradation, which further impacts signal-to-interference-plus-noise ratio (SINR) and spectral efficiency (SE). The newly reported angular spreads in [7] implies that modifications are required for beam management in the new goldilocks FR3 mid-band spectrum. For sensing applications in ISAC, this also means that precise and accurate beam alignment is needed for target detection, localization and tracking. In addition to channel modeling, the call for new metrics and figures-of-merit, such as the Waste Factor, offer potential for energy efficiency optimization in $6$G wireless systems.
II-B FR3 frequency-dependent features and losses
As discussed in the previous section, empirical field data [8] in InF (dense-scatterer environment) indicate frequency-dependent trends for DS and angular spread (AS). This is not suprising, as frequency-dependence propagation features were measured and revealed in [13] for both outdoors and indoors empirically discovered in what eventually became the FR3 bands of $28$ and $73\operatorname{GHz}$. For modeling the channel path loss over distance, extensive empirical measurements show the close-in free space model with a frequency-weighted path loss exponent (CIF) is more suited for indoors [13], and extends the CI model which has a physics-based close-in free space reference distance as a leverage point for the slope of the exponential path loss, while incorporating the frequency dependence feature of path loss observed indoors. The multipath temporal and spatial statistics at the lower part of the FR3, e.g. $7\operatorname{GHz}$, varies considerably compared to the higher part of FR3, e.g. $16$-$24\operatorname{GHz}$, which confirms less multipath dispersion in both time and space at higher mmWave frequencies [1]. Perhaps most importantly, rain attenuation and foliage losses vary significantly across the FR3. Following ITU-R recommendation (P.838838838838-$3$), the specific attenuation at $7\operatorname{GHz}$ for heavy rain (corresponds to rain rate of $8\operatorname{mm/hr}$) is $0.04\operatorname{dB/km}$, whereas at $24\operatorname{GHz}$, the specific attenuation is $1.16\operatorname{dB/km}$, i.e. a gap of $1.16\operatorname{dB}$ difference over $1\operatorname{km}$. Following Weissberger’s model, the foliage loss difference at $24\operatorname{GHz}$ relative to that at $7\operatorname{GHz}$ is $14.52\operatorname{dB}$ at $100\operatorname{m}$. In particular, the foliage loss is $34.66\operatorname{dB}$ at $7\operatorname{GHz}$ over $100\operatorname{m}$, whereas it is $49.18\operatorname{dB}$ at $24\operatorname{GHz}$ over $100\operatorname{m}$. However, the link can be improved to extend the transmitter’s coverage area via coherent multi-beam combining [12].
III Spectrum Agility & Coverage-Rate Tradeoffs
III-A Spectrum Agility for Coverage-Rate Tradeoffs
The lower FR3 frequencies provide better coverage, as detailed in Section II, hence a fundamental question arises: What mechanism can optimally balance rate-coverage tradeoffs ? A particularly interesting ray tracing simulation was conducted in [14] to study the coverage and rate statistical behaviour across the FR3 spectrum. In terms of rate, with no blockages, and as expected, it was shown that $24\operatorname{GHz}$ usage dominates in terms of rate for outdoors. Interestingly, when a 3GPP blockage model B was used, it was shown that there was a fallback need towards $12\operatorname{GHz}$ and $6\operatorname{GHz}$ for maximum rate. But even more, and making use of the the comprehensive and detailed penetration models and findings from NYU WIRELESS [1], it is observed that while materials like glass and wood exhibit relatively permeable characteristics at FR3 frequencies, others, e.g. low-emissivity glass, are not permeable. A particularly intriguing observation is that concrete, which is relatively permeable at sub-$6$ $\operatorname{GHz}$ frequencies, becomes completely impermeable at FR3. Additionally, as reported in [14], the indoor user rate distribution indicates that users operating at $24\operatorname{GHz}$ will need to fallback to lower frequencies, which means that, if we really want to leverage the large FR3 spectrum, frequency hopping mechanisms between FR3 bands should be designed to account for the various blockages and penetration losses. The diversity and hopping mechanism naturally pushes forward the need for a spectrum agility in the FR3. Besides, the agility also inherently secures certain transmissions from malicious users and eavesdroppers. A natural question then to ask would be what hopping mechanism should be adopted between the FR3 bands in order to be spectrally efficient and agile?
III-B Incumbents: Coexistence & Opportunity
The primary challenge of utilizing FR3 for $6$G lies in regulatory issues. The spectrum is densely populated with both civilian and government incumbents for various applications beyond fixed and mobile wireless access, like meteorology, radio astronomy, and maritime radio navigation. In contrast to $5$G, one would need to pay special care towards such incumbents. Even if regulators reach consensus on spectrum availability and licensing frameworks, the most technically demanding aspect will be determining how to share this spectrum without interfering with existing users. Note that satellite users are already spectrum-sharing amongst themselves. Vice-versa, we can also consider a fundamental question that is: are commercial services and malicious users able to harm satellites? A possible approach would be through interference nulling to guarantee protection of satellite uplink (UL) communications without sacrificing much the terrestrial downlink (DL), which translates into a greater chance of co-existence. Sharing spectrum efficiently while avoiding interference, especially in cases where users operate close to one another, is a technical challenge that requires effective coordination mechanisms. Relocating incumbent users can be expensive, often eclipsing the auction value of the spectrum itself. Inspired by spectrum refarming mechanisms, dynamic spectrum sharing (DSS) can ensure backward compatibility with older generations in a dynamic way. More specifically, DSS allows service providers to bypass a static cut between different generations in frequency, allowing for wiser spectral utilization. Similarly, DSS can allow a dynamic smart way of sharing spectrum between $6$G users and incumbents within the band, but this would be very challenging.
IV Massive MIMO, Compact Arrays & Power Considerations
Doubling the frequency implies a loss by a factor of $4$ in Watts. However, this loss can be regained by filling the same effective area, covering the space, which also means more compact (but not bigger) arrays, which is referred to as antenna gain. Given half a wavelength spacing, then multiplying the carrier frequency by $N$ necessitates $N^{2}$ more antennas to cover the aperture area. For FR3, transitioning from $3.5\operatorname{GHz}$ to $7\operatorname{GHz}$ faces a $6\operatorname{dB}$ free space path-loss, but also a $6\operatorname{dB}$ antenna gain assuming the same aperture size, enabling similar coverage to lower frequencies with higher efficiency. Meanwhile, the coherence distance shrinks with higher frequencies, which implies that the minimum distance over which signals are uncorrelated in phase and amplitude also decreases, giving rise to the need of packing up more antennas in a given aperture. The challenge remains in the NLoS case, especially for neutral host, where there are much more severe penetration losses at higher frequencies compared to lower counterparts, as mentioned in Sections II and III, in addition to diffraction losses which scale as $10\log f$.
[…]
It is worth noting that new approaches have emerged for the design of ultra-wide band (UWB) antenna arrays whereby mutual coupling (MC) is intentionally introduced between different elements to expand the bandwidth. As far as massive MIMO (mMIMO) is concerned, it was shown that compact mMIMO can open extra degrees-of-freedom in frequency domain [15]. Specifically, MC between array elements can be leveraged in tightly coupled colinear antenna arrays to expand the bandwidth. While interesting to attain higher rates, how can sensing be performed under intentionally induced MC in compact mMIMO designs, especially given its frequency dependence across the UWB?
IV-B Indoor reflection modeling for extended targets
The reflections arising from a single target share a common relation simply because they arise from the same target, i.e. extended target (ET). ETs can reveal information besides location, e.g. size, shape, and orientation, which can enable use cases like gesture recognition. In fact, ET modeling is particularly relevant to indoors owing to its property of rich multipath reflections. Therefore, more modeling efforts should be conducted to characterize ET on lower and higher counterparts of FR3. Even more, as alluded to in Section IV-A, and besides UWB, mMIMO can allow resolving ET paths in spatial domain.
IV-C mMIMO architectures: Fully-digital vs. hybrid
As number of antennas grow large to support various FR3 ISAC applications, which will heavily depend on SE and multi-band operations, a fully-digital scheme may seem complex, due to high power consumption issues. For ISAC, one can consider a fully digital array employing low-resolution digital-to-analog and analog-to-digital converters to trade-off quantization noise with power consumption. Therefore, we can ask what is an optimal fully digital design that can balance sensing tasks with reasonable power consumption ? In contrast, hybrid beamforming seems like a good tradeoff to alleviate the power consumption problem, by trading-off some beamforming and beam selection gains, which, in turn has negative reverberation effects on the received SINR and SE for communications, in addition to detection and localization accuracy for sensing. Therefore, what is the optimal analog beamforming number of sub-panels to strike such trade-offs in hybrid design for sensing and communication (S&C)?
IV-D Coherence time across FR3
Higher frequencies naturally reduce the coherence time, which in turn increases the complexity of channel estimation and makes spatial multiplexing more challenging. An inherent property of FR3 is that the coherence time can significantly vary across FR3, especially in mobile and highly dynamic environments. The coherence block shrinks at higher frequencies, hence the channel outdates faster. In turn, this can limit the duration available for multiplexing data, which can also limit the number of terminals that can be multiplexed simultaneously due to constraints on orthogonal pilot sequences fitting within the coherence time. Within the coherence time, we also know that the achievable rate grows as $\min\left(N_{T},N_{R}\right)\log_{2}(1+\operatorname{SNR})$ per unit of frequency. Hence, mMIMO can allow for more data multiplexing in an environment where less coherence time is available, especially at the higher end of FR3 bands. However, special attention is needed for channel state information acquisition. A question here could be: Should higher FR3 frequencies be dedicated to less mobile environments, while lower FR3 bands to highly dynamic environments ?
V Multiband sensing for FR-3 ISAC
ISAC use cases planned for FR3 mean that channel models must support S&C functions. Multi-band sensing for future systems can be realized through non-contiguous aggregation, whereby different frequency blocks are spread across different bands or parts of the FR3 spectrum.
Next, we provide an insightful analysis of some aspects of multiband sensing. As frequency increases, there is typically more spectrum available, and more antennas can be packed per unit of space. Thus, we assume that the bandwidth and antenna gain is proportional to the center frequency. It can be verified that when the frequency is doubled, the signal-to-noise ratio (SNR) and the Cramér-Rao bound (CRB) drop by a factor of $2$. An example is provided in Fig. 2. The drop of SNR at higher frequencies shifts the waterfall region, and the lower frequencies have more margin to accurately start estimating the delay as predicted by the CRB. Conversely, since the CRB of the higher frequency system is smaller, it means that after the waterfall region, the higher frequency achieves a more accurate estimate.
Figure 2: Multi-frequency ISAC trade-offs between low and high frequencies.
Our analysis reveals two insights for leveraging multi-band sensing. The first is to obtain estimates individually and combine via a maximum ratio combining type approach. Specifically, one can assign more weights to estimates obtained at higher frequencies. The second is relevant for bistatic sensing, where we can rely on higher FR3 frequencies to obtain synchronization via LoS path due to the need of accurate timing, but then we can tolerate to fallback to lower frequencies, which we know can cover more targets. Based on this scheme, a good time difference of arrival location-based method can be developed leveraging the entire multi-band information on FR3.
VI Key FR3 features in $6$G Standards
This section explores key modifications envisioned for next-generation $6$G cellular standardization to unlock the full potential of spectrally agile FR3 radios. Specifically, we examine both 3GPP innovations—focusing on radio frame enhancements tailored for spectral agility—and new O-RAN-based features, particularly in O-RAN open radio unit (O-RU) designs, that facilitate intelligent spectrum utilization.
3GPP
The challenges posed by FR3, e.g. wider bandwidths, non-contiguous carriers, increased propagation losses, and dynamic spectrum environments, necessitate innovations in frame design to ensure low-latency, high-reliability, and efficient multi-band operations. Below, we outline key advancements required to enhance spectral agility, synchronization, and resource management in next-generation wireless systems:
  • Multi-band aggregation, i.e. combining separate, non-contiguous spectrum blocks to optimize the balance between performance and complexity, enhancing both S&C capabilities.
  • Dynamic slot configuration for multi-band coordination, which implements AI-driven adaptive slot structures that dynamically adjust frame configurations across multiple bands. The dynamic strategy ensures optimized latency, throughput, and reliability by tailoring subframe allocations to each band’s propagation characteristics.
  • Multi-band pilot design, which aims at designing hybrid pilot structures that combine common and band-specific pilots while employing sparse allocation, frequency sharing, and adaptive density. Multi-Band pilot design can be realized via AI-driven optimization and compressed sensing techniques to minimize redundancy and intelligently manage pilots for greater efficiency in multi-band environments.
  • Agile timing advance, which aims at developing new timing advance strategies to ensure seamless switching and synchronization across carriers in different frequency bands, which is critical for achieving low-latency and high-reliability performance in $6$G.
  • Multi-carrier synchronization and scalable numerology, meaning that it is vital to introduce enhanced scalable numerology to support multi-carrier synchronization across diverse FR3 bandwidths, which further allows adaptive subcarrier spacing and multi-band coordination without excessive signaling overhead.
[…]
From RAN perspective, the open radio access network (O-RAN) RAN intelligent controller (RIC) plays a crucial role in enabling spectral agility and intelligence by leveraging real-time (RT) and non-RT network optimizations. Through the use of xApps (near-RT RIC applications) and rApps (Non-RT RIC applications), the RIC can dynamically manage spectrum resources, adapt to changing conditions, and optimize performance in FR3. Hereafter, we examine how spectral agility and spectral intelligence can be achieved through the RIC.
Spectral Agility: The ability of a network to dynamically reconfigure spectrum usage based on real-time conditions, interference levels, traffic demands, and regulatory constraints can be achieved through the following mechanisms:
  • Dynamic carrier aggregation & multi-band adaptation: The Near-RT RIC can monitor real-time spectrum conditions and trigger carrier aggregation adjustments across multiple FR3 bands, which enables on-the-fly switching between licensed, shared, and unlicensed spectrum resources, optimizing throughput and reducing congestion.
  • Intelligent spectrum sensing and sharing: By integrating AI/ML models within rApps, the Non-RT RIC can analyze historical spectrum usage patterns and proactively allocate spectrum to different cells or operators, thus enabling dynamic spectrum access and coexistence with satellite, radar, or other wireless services operating in FR3.
Spectral Intelligence: Efficiency, reliability, and adaptability of spectrum can be achieved through intelligent decision-making and automation.
  • Beamforming, channel estimation, and interference management: xApps in the Near-RT RIC continuously monitor spectrum conditions and adjust beamforming strategies for multi-user MIMO in FR3. Advanced channel estimation and interference prediction can help minimize co-channel and adjacent-band interference, ensuring optimal signal quality.
  • Multi-band O-RU control for distributed MIMO: The Near-RT RIC can coordinate multi-band O-RUs across FR3 to enable distributed MIMO architectures. Leveraging O-RAN’s open fronthaul interface, it ensures tight synchronization across spatially separated RUs, maximizing spectral efficiency.
  • Power efficiency: Other key innovations that can be implemented by leveraging the RIC are adaptive sampling and dynamic range control. By adjusting these parameters based on real-time spectral conditions, FR3 O-RUs reduce unnecessary power consumption and enhance operational efficiency.
VII Conclusions
This paper outlined a vision for $6$G in the FR3 spectrum, exploring its upper mid-band potential, key challenges, and necessary standardization changes. We highlighted FR3’s advantages in capacity, coverage, and SE, while emphasizing the impact of its channel characteristics on beam management. Based on our FR3 measurements, we underscored the need for frequency hopping to mitigate blockages, maximize SE, and incumbent coexistence, reinforcing the importance of spectrum agility. We examined the potential of mMIMO for FR3 and discussed design challenges, particularly in the uplink. For sensing, we analyzed performance bounds for FR3 ISAC multiband sensing. Additionally, we detailed key $6$G standardization features, including 3GPP radio frame modifications and O-RAN-based innovations, using RIC-driven xApps and rApps for intelligent spectrum management. However, open questions remain regarding FR3 propagation characteristics, PHY design, and regulatory considerations.

### v2

@@
 I-A The FR3 Spectrum
-As the global wireless industry advances towards $6$G, the upper mid-band frequencies, being a part of the high super high frequency (SHF) bands, are critical resources due to their balance between coverage and bandwidth availability, compared to lower frequency range 1 (FR1) (up to $7.125\operatorname{GHz}$ [3]) and higher frequency range 2 (FR2) ranges, which includes $24.25\operatorname{GHz}$ to $71\operatorname{GHz}$ [4].
+As the global wireless industry advances towards $6$G, the upper mid-band frequencies, being a part of the high super high frequency (SHF) bands, are critical resources due to their balance between coverage and bandwidth availability, compared to lower frequency range 1 (FR1) (up to $7.125\operatorname{GHz}$) and higher frequency range 2 (FR2) ranges, which includes $24.25\operatorname{GHz}$ to $71\operatorname{GHz}$ [3].
 Often termed the “Golden Band”, or ”Goldilocks Spectrum”, for $6$G, FR3 frequencies ($7.125$-$24.25$ $\operatorname{GHz}$ [2] as depicted in Fig.
@@
 The FR3 spectrum holds growing importance and interest for industry and regulatory bodies like the U.S.
-National Telecommunications and Information Administration (NTIA), the $3^{\text{rd}}$ Generation Partnership Program (3GPP) [5] and the Federal Communications Commission (FCC) , who are currently evaluating its potential alongside existing mobile radio bands in order to expand cellular services.
-Besides mobile operator use, the upper mid-band is being considered to coexist with incumbent services such as satellite communications, radio astronomy, and earth exploration, in addition to warfare activities, such as the next generation jammer mid-band (NGJ-MB) operating within $509\operatorname{MHz}$ to $18\operatorname{GHz}$ band [6].
+National Telecommunications and Information Administration (NTIA), the $3^{\text{rd}}$ Generation Partnership Program (3GPP) and the Federal Communications Commission (FCC) , who are currently evaluating its potential alongside existing mobile radio bands in order to expand cellular services.
+Besides mobile operator use, the upper mid-band is being considered to coexist with incumbent services such as satellite communications, radio astronomy, and earth exploration, in addition to warfare activities, such as the next generation jammer mid-band (NGJ-MB) operating within $509\operatorname{MHz}$ to $18\operatorname{GHz}$ band.
 Figure 1: The spectrum showing the placement of FR3 Golden band, relative to its other counterparts according to different bodies and in various regions [1].
@@
 16.95 11–97 1.32 2.66 3.05 8.11 22.1 40.7 34.2 43.5
-28 4–46 1.2 1.8 2.7 9.7 10.8 17.1 Med: 39.1 Med: 31.8
+28 4–46 1.2 [6] 1.8 2.7 [6] 9.7 10.8 17.1 Med: 39.1 Med: 31.8
 UMi [4] 6.75 40–1000 1.79 2.57 2.56 6.53 62.8 75.6 16.79 25.62
 16.95 40–1000 1.85 4.05 2.59 8.78 46.5 65.8 18.23 19.08
-28 31–187 2.02 8.98 3.56 8.91 26.7 46.3 Med:14 Med:30
+28 31–187 2.02 - 2.1 [7] 8.98 3.4-3.56[7] 8.91 26.7 46.3 Med:14 Med:30
 InF [5] 6.75 9–38 1.39 1.86 1.78 2.46 14 30.3 23.71 66.38
@@
 The findings of FR3 channel measurements conducted by NYU WIRELESS in [1, 4, 5] reveal that for wideband channels at $16.95\operatorname{GHz}$ in the FR3 band, the omnidirectional pathloss exponent (PLE) values, synthesized from directional channel measurements, in LoS scenarios were slightly lower than those in milli-meter wave (mmWave), indicating less signal attenuation and more of a waveguide effect over distance, e.g.
-the omnidirectional LoS PLEs in InH, UMi and InF for $6.75\operatorname{GHz}$, $16.75\operatorname{GHz}$, and $28\operatorname{GHz}$ are given in Table I, e.g.
-$1.34$ at $6.75\operatorname{GHz}$, $1.32$ at $16.75\operatorname{GHz}$, and $1.2$ at $28\operatorname{GHz}$ [1].
-In UMi, the omnidirectional LoS PLEs are $1.79$ at $7\operatorname{GHz}$ [1], $1.85$ at $16.75\operatorname{GHz}$ [1], and $2.02$ at $28\operatorname{GHz}$ [1], which indicates slightly less loss over distance than free space channel (eg when PLE $=2$) and with less loss at lower frequencies.
+the omnidirectional LoS PLEs in InH, UMi and InF for $6.75\operatorname{GHz}$, $16.95\operatorname{GHz}$, and $28\operatorname{GHz}$ are given in Table I, e.g.
+$1.34$ at $6.75\operatorname{GHz}$, $1.32$ at $16.95\operatorname{GHz}$, and $1.2$ at $28\operatorname{GHz}$ [1].
+In UMi, the omnidirectional LoS PLEs are $1.79$ at $6.75\operatorname{GHz}$ [1], $1.85$ at $16.95\operatorname{GHz}$ [1], and $2.02$ at $28\operatorname{GHz}$[7], which indicates slightly less loss over distance than free space channel (e.g., when PLE $=2$) and with less loss at lower frequencies.
 Although antenna patterns and gains vary widely, and higher frequencies enable the use of directional antennas that can offset channel loss in mmWave bands, 3GPP and researchers often use an omnidirectional channel model to standardize link analysis and ensure consistency in evaluations while enabling the use of any type of directional pattern to be applied to the models [9], [10].
 NLoS PLEs were found to be higher than LoS scenario as is found in all 3GPP bands as per Table I, e.g.
-NLoS PLE for UMi of $2.56$ at $6.75\operatorname{GHz}$ [1], $2.59$ at $16.75\operatorname{GHz}$ [1], and for InF the NLoS PLE is $1.78$ at $6.75\operatorname{GHz}$ [8], and $2.11$ at $16.75\operatorname{GHz}$ [8], the loss over distance is much lower (eg $14.5\operatorname{dB}$ per decade less) than those observed at mmWave frequencies, e.g.
-NLoS PLE of $3.56$ at $28\operatorname{GHz}$, highlighting how FR3 offers improved coverage for a comparable equivalent isotropic radiated power (EIRP) and radio frequency (RF) bandwidth relative to higher frequencies.
-Referring to Table I, the PLE for LoS in UMi is $1.79$ for $6\operatorname{GHz}$ and $2.02$ for $28\operatorname{GHz}$, implying greater coverage with a $2.2\operatorname{dB}$ stronger signal per decade of distance, resulting in, $4.4\operatorname{dB}$ over $100\operatorname{m}$ and $6.6\operatorname{dB}$ stronger signal over $1\operatorname{km}$, which clearly demonstrates how a lower PLE translates into significant coverage improvement, which may be traded for higher capacity for users within a cell, a factor that will be of paramount importance to carriers when deploying $6$G technologies.
+NLoS PLE for UMi of $2.56$ at $6.75\operatorname{GHz}$ [1], $2.59$ at $16.95\operatorname{GHz}$ [1], and for InF the NLoS PLE is $1.78$ at $6.75\operatorname{GHz}$ [5], and $2.11$ at $16.95\operatorname{GHz}$ [5], the loss over distance in NLoS locations is lower (e.g., $12$ to $14.5\operatorname{dB}$ per decade less) than those observed at mmWave frequencies, e.g.
+NLoS PLE of $3.4$ to $3.56$ at $28\operatorname{GHz}$ UMi [7], highlighting how FR3 offers improved coverage in NLOS channels relative to higher-frequency mmWave bands.
+It is important to note, however, that practical systems use higher antenna gains at higher frequencies, and a common error is to attempt to use equivalent isotropic radiated power (EIRP) to compare coverage at different frequencies, since the transmit and receive antenna physical apertures are proportionally reduced with shrinking wavelength if antenna gains are left constant over frequency, such that equal power transmitters with EIRP over different frequency will give the false impression that higher bands have far less coverage.
+In practice, the proper comparison is to consider equal antenna aperture areas at different frequencies, where in practice the extra array gain at mmWave can overcome the additional channel loss by many dB [11].
+Referring to Table I, the PLE for LoS in UMi is $1.79$ for $6.75\operatorname{GHz}$ and $2.02$ to $2.1$ [7] for $28\operatorname{GHz}$, implying greater coverage with a $2.2\operatorname{dB}$ stronger signal per decade of distance, resulting in, $4.4\operatorname{dB}$ over $100\operatorname{m}$ and $6.6\operatorname{dB}$ stronger signal over $1\operatorname{km}$, which clearly demonstrates how a lower PLE translates into significant coverage improvement, which may be traded for higher capacity for users within a cell, a factor that will be of paramount importance to carriers when deploying $6$G technologies.
 The time delay spread of a propagation channel has impact on signaling formats, in addition to accuracy for integrated sensing and communications (ISAC) applications such as position location and multi-user synchronization methods.
@@
 Furthermore, the measurements in [1] found that higher FR3 frequencies, i.e.
-$16.75\operatorname{GHz}$, exhibited a narrower RMS angular spread compared to the higher FR1 frequency at $6.75\operatorname{GHz}$, hence indicating fewer, more focused multipath components, which is beneficial for spatial multiplexing as it reduces interference between signal paths and allows more precise beamforming.
+$16.95\operatorname{GHz}$, exhibited a narrower RMS angular spread compared to the higher FR1 frequency at $6.75\operatorname{GHz}$, hence indicating fewer, more focused multipath components, which is beneficial for spatial multiplexing as it reduces interference between signal paths and allows more precise beamforming.
 Regarding material penetration, losses were consistently higher at $16.95\operatorname{GHz}$ in FR3 than at lower frequencies, with losses dependent on material type and polarization configuration, confirming the inherent limitations of FR3 in penetrating certain materials, such as low-emissivity (IRR) glass and concrete [1], yet able to allow more penetration that at FR3 mmWave [1].
@@
 III-A Spectrum Agility for Coverage-Rate Tradeoffs
-The lower FR3 frequencies provide better coverage, as detailed in Section II, hence a fundamental question arises: What mechanism can optimally balance rate-coverage tradeoffs ?
+Under a constant EIRP with shrinking aperture assumption, the lower FR3 frequencies provide better coverage, as detailed in Section II.
+If instead the antenna physical area is held fixed, the extra array gain available at mmWave more than compensates for the additional free space pathloss.
+A fundamental question arises: What mechanism can optimally balance rate-coverage tradeoffs ?
 A particularly interesting ray tracing simulation was conducted in [14] to study the coverage and rate statistical behaviour across the FR3 spectrum.
@@
 A question here could be: Should higher FR3 frequencies be dedicated to less mobile environments, while lower FR3 bands to highly dynamic environments ?
+IV-E RF frontend efficiency and array integration limits
+Energy efficiency of power amplifiers can becoem a challenge in the FR3 band, so transmit side losses can limit the link budget advantages.
+Particularly, packing large numbers of elements into a fixed aperture soon hits practical barriers due to thermal dissipation, EIRP limits, and fronthaul capacity.
+Therefore, from array design and RF frontend perspective, how can future FR3 radios achieve high energy efficiency in their power amplifiers while packing dense antenna arrays under real world thermal, EIRP, and processing limits, without giving up the very coverage-rate gains the band is meant to provide?
+Waste factor can serve as a fundamental metric towards the design.
 V Multiband sensing for FR-3 ISAC

### v3

@@
 spectral efficiency
+PN
+phase noise
 AR
@@
 mmWave
-milli-meter wave
+millimeter wave
 gNB
@@
 near-field
-FF
+$F$
 far-field
@@
 angular spread of arrival
-Upper Mid-Band Spectrum for 6G: Vision, Opportunity and Challenges
-Ahmad Bazzi, Roberto Bomfin, Marco Mezzavilla, Sundeep Rangan, Theodore Rappaport, Marwa Chafii Ahmad Bazzi, Roberto Bomfin, and Marwa Chafii are with the Engineering Division, New York University Abu Dhabi (NYUAD), 129188, UAE (email: {ahmad.bazzi,roberto.bomfin, marwa.chafii}@nyu.edu)Ahmad Bazzi, Marwa Chafii, Sundeep Rangan, and Theodore Rappaport are with NYU WIRELESS, NYU Tandon School of Engineering, Brooklyn, 11201, NY, USA (email: {srangan,tsr}@nyu.edu).Marco Mezzavilla is with the Telecommunications Engineering at the Department of Electronics, Information, and Bioengineering (DEIB) at Politecnico di Milano, Milan, Italy.
+IEEE
+Institute of Electrical and Electronics Engineers
+IMT
+International Mobile Telecommunication
+CRLB
+Cramér-Rao lower bound
+RSS
+received signal strength
+DMC
+dense multipath components
+$E$
+energy efficiency
+Upper Mid-Band Spectrum for 6G:
+Vision, Opportunity and Challenges
+Ahmad Bazzi, Roberto Bomfin, Marco Mezzavilla, Sundeep Rangan, Theodore S.
+Rappaport, Marwa Chafii Ahmad Bazzi, Roberto Bomfin, and Marwa Chafii are with the Engineering Division, New York University Abu Dhabi (NYUAD), 129188, UAE (email: {ahmad.bazzi,roberto.bomfin, marwa.chafii}@nyu.edu)Ahmad Bazzi, Marwa Chafii, Sundeep Rangan, and Theodore Rappaport are with NYU WIRELESS, NYU Tandon School of Engineering, Brooklyn, 11201, NY, USA (email: {srangan,tsr}@nyu.edu).Marco Mezzavilla is with the Telecommunications Engineering at the Department of Electronics, Information, and Bioengineering (DEIB) at Politecnico di Milano, Milan, Italy.
 (email: marco.mezzavilla@polimi.it).
@@
 In this paper, we chart a compelling vision for 6G within the frequency range 3 (FR3) spectrum, i.e.
-$7.125$-$24.25$ $\operatorname{GHz}$, by delving into its key enablers and addressing the multifaceted challenges that lie ahead for these new frequency bands.
-Here we highlight the physical properties of this never-before used spectrum by reviewing recent channel measurements for outdoor and indoor environments, including path loss, delay and angular spreads, and material penetration loss, all which offer insights that underpin future 5G/6G wireless communication designs.
-Building on the fundamental knowledge of the channel properties, we explore FR3 spectrum agility strategies that balance coverage and capacity (e.g.
-data rate) tradeoffs, while also examining coexistence with incumbent systems, such as satellites, radio astronomy, and earth exploration.
-Moreover, we discuss the potential of massive multiple-input multiple-output, compact and digital architectures, and evaluate the potential of multiband sensing for FR3 integrated sensing and communications.
+7.1257.125-$24.25$ $\operatorname{GHz}$, by delving into its key enablers and addressing the multifaceted challenges that lie ahead for these new frequency bands.
+Here we highlight the physical properties of this never-before used spectrum for cellular by reviewing recent channel measurements for outdoor and indoor environments, including path loss, delay and angular spreads, and material penetration loss, all which offer insights that underpin future 5G/6G wireless communication designs.
+Building on the fundamental knowledge of the channel properties, we explore FR3 spectrum agility strategies that balance coverage and capacity tradeoffs, while examining coexistence with incumbent systems, such as satellites, radio astronomy, and earth exploration.
+Moreover, we discuss the potential of massive multiple-input multiple-output technologies, challenges for commercial deployment, and potential solutions for FR3, including multiband sensing for FR3 integrated sensing and communications.
 Finally, we outline 6G standardization features that are likely to emerge from 3GPP radio frame innovations and open radio access network developments.
 Index Terms:
-$6$G, upper midband, FR3, channel characteristics, spectrum agility, ISAC
+66G, upper midband, FR3, channel characteristics, spectrum agility, ISAC
 I Introduction
-By $2034$, global mobile data traffic is expected to grow by $5$x–$9$x, with artificial intelligence (AI) accounting for one-third of the traffic.
-There is an industry-wide consensus that $6$G is expected to launch around $2030$ with new spectrum in frequency range 3 (FR3) [1], which spans $7.125$-$24.25$ $\operatorname{GHz}$ [2].
-To meet the $6$G timeline, it is crucial to allocate the necessary spectrum a few years in advance, in order to ensure incumbents are accommodated and that the technology is ready for mass adoption at the global launch of $6$G.
+By $2034$, global mobile data traffic is expected to grow by five- to nine-fold, with artificial intelligence (AI) accounting for one-third of the traffic.
+There is an industry-wide consensus that 66G is expected to launch around $2030$ with new spectrum in frequency range 3 (FR3) [1], which spans 7.1257.125-$24.25$ $\operatorname{GHz}$ [2].
+To meet the 66G timeline, it is crucial to allocate the necessary spectrum a few years in advance, in order to ensure incumbents are accommodated and that the technology is ready for mass adoption at the global launch of 66G.
 Companies are actively conducting proof-of-concept product trials to address bottlenecks and tackle myriad implementation challenges.
-At the same time, regulators must determine how the spectrum will be allocated, and operators need to develop viable business cases and deployment/integration strategies for nationwide $6$G networks.
+At the same time, regulators must determine how the spectrum will be allocated, and operators need to develop viable business cases and deployment/integration strategies for nationwide 66G networks.
 I-A The FR3 Spectrum
-As the global wireless industry advances towards $6$G, the upper mid-band frequencies, being a part of the high super high frequency (SHF) bands, are critical resources due to their balance between coverage and bandwidth availability, compared to lower frequency range 1 (FR1) (up to $7.125\operatorname{GHz}$) and higher frequency range 2 (FR2) ranges, which includes $24.25\operatorname{GHz}$ to $71\operatorname{GHz}$ [3].
-Often termed the “Golden Band”, or ”Goldilocks Spectrum”, for $6$G, FR3 frequencies ($7.125$-$24.25$ $\operatorname{GHz}$ [2] as depicted in Fig.
+As the global wireless industry advances towards 66G, the upper mid-band frequencies, being a part of the high super high frequency (SHF) bands, are critical resources due to their balance between coverage and bandwidth availability, compared to lower frequency range 1 (FR1) (up to $7.125\operatorname{GHz}$) and higher frequency range 2 (FR2) ranges, which includes $24.25\operatorname{GHz}$ to $71\operatorname{GHz}$ [3].
+Often termed the “Golden Band”, or ”Goldilocks Spectrum”, for 66G, FR3 frequencies (7.1257.125-$24.25$ $\operatorname{GHz}$ [2] as depicted in Fig.
 1 [1]) are particularly suited for enhancing network capacity while maintaining reasonable propagation characteristics, offering moderate propagation losses, enabling extensive urban and suburban reach using existing towers.
+U.S.
+regulators have signalled that 7.257.25-$7.4$ GHz will serve as an initial slice for FR3, which is favorable as it avoids the $22$ GHz water-vapour and $60$ GHz oxygen absorption peaks, keeping rain attenuation modest enough for macro-cell coverage.
+Further FR3 spectrum is expected to be layered outward from this starter band.
 The FR3 spectrum holds growing importance and interest for industry and regulatory bodies like the U.S.
 National Telecommunications and Information Administration (NTIA), the $3^{\text{rd}}$ Generation Partnership Program (3GPP) and the Federal Communications Commission (FCC) , who are currently evaluating its potential alongside existing mobile radio bands in order to expand cellular services.
-Besides mobile operator use, the upper mid-band is being considered to coexist with incumbent services such as satellite communications, radio astronomy, and earth exploration, in addition to warfare activities, such as the next generation jammer mid-band (NGJ-MB) operating within $509\operatorname{MHz}$ to $18\operatorname{GHz}$ band.
+Besides mobile operator use, the upper mid-band is being considered to coexist with incumbent services such as satellite communications, radio astronomy, and earth exploration, in addition to warfare activities, such as the next generation jammer mid-band operating within $509\operatorname{MHz}$ to $18\operatorname{GHz}$ band.
 Figure 1: The spectrum showing the placement of FR3 Golden band, relative to its other counterparts according to different bodies and in various regions [1].
 Notably, the NTIA highlighted specific FR3 bands, including $7.125\operatorname{GHz}$-$8.4\operatorname{GHz}$, and $14.8$-$15.35\operatorname{GHz}$, as part of the International Telecommunication Union (ITU) World Radio Conference $2023$ (WRC-$23$) agenda, emphasizing the pending authorization in the USA and hence the need for precise channel modeling to ensure effective spectrum utilization and sharing.
+I-B Improvement items by ITU, 3GPP & Organization
+3GPP work started in 2024 during Release 19, including channel models for the FR3.
+The recent FR3 measurements by NYU WIRELESS are unified in Section II.
+Meanwhile, the ITU has identified a list of capabilities in International Mobile Telecommunication (IMT)-$2030$ in the form of improvements and additional capabilities compared to those in IMT-$2020$.
+The improvements include: a peak data rate of up to $200\operatorname{Gbps}$, which is $10\times$ that of IMT-$2020$, in addition to user experienced data rates, which are $300\operatorname{Mbps}$ and $500\operatorname{Mbps}$, i.e a factor of $3$ to $5$ improvement.
+The area traffic capacity are expected to be $30$ and $50\operatorname{Mbps}$ per square meter and the spectrum efficiency was set to an improvement of $1.5$ and $3$ higher than that of IMT-$2020$, while connection density is set to a target of $10^{6}$-$10^{8}$ devices per km2.
+In Section III, we discuss different methods to achieve the aforementioned factors.
+Moreover, we highlight related challenges for commercial deployment in Section IV.
+Moreover, IMT-2030 also calls for enhancing radio access network (RAN), including flexibility, intelligence, and resiliency, so that a single RAN can be sliced, AI-optimized, and seamlessly span terrestrial and non-terrestrial links while guaranteeing specific QoS.
+Besides RAN, we mention several open questions and provide potential solutions in Section V.
 II FR3 Channel Characteristics
 TABLE I: Omni PLEs from the CI PL model with 1 m reference distance, Omni DS, and Omni ASA at RX [1, 4, 5].
-Env Freq. (GHz) Dist (m) LOS NLOS LOS Omni RMS DS NLOS Omni RMS DS LOS Omni RMS ASA NLOS Omni RMS ASA
-Omni PLE Sigma Omni PLE Sigma
-InH [1] 6.75 11–97 1.34 3.51 2.72 9.21 37.7 48 40.9 58.2
-16.95 11–97 1.32 2.66 3.05 8.11 22.1 40.7 34.2 43.5
-28 4–46 1.2 [6] 1.8 2.7 [6] 9.7 10.8 17.1 Med: 39.1 Med: 31.8
-UMi [4] 6.75 40–1000 1.79 2.57 2.56 6.53 62.8 75.6 16.79 25.62
-16.95 40–1000 1.85 4.05 2.59 8.78 46.5 65.8 18.23 19.08
-28 31–187 2.02 - 2.1 [7] 8.98 3.4-3.56[7] 8.91 26.7 46.3 Med:14 Med:30
-InF [5] 6.75 9–38 1.39 1.86 1.78 2.46 14 30.3 23.71 66.38
-16.95 9–38 1.75 3.1 2.11 3.29 12.7 29 14.32 50.4
 II-A FR3 Channel Modeling
@@
 Integrated channel models are indispensable for addressing diverse propagation scenarios, ranging from urban environments to free space, yet 3GPP developed the TR 38.901 channel models without many field measurements across much of the $100\operatorname{GHz}$ wide swath of the spectrum, and without any measurements from the FR3 bands.
-As a result, the model is an estimate, formulated from sparse measurements across particular bands within the $0.5$ to $100\operatorname{GHz}$ range.
+Consequently, the model is an estimate, formulated from sparse measurements across particular bands within the $0.5$ to $100\operatorname{GHz}$ range.
 Channel modeling research for spectra above $6\operatorname{GHz}$ was completed in Release 17 of TR $38.901$ in April $2022$, offering a comprehensive set of models for evaluating various physical layer technologies.
-However, the primary focus of $5$G channel modeling has been on frequencies below $6\operatorname{GHz}$ (FR1) and above $24\operatorname{GHz}$ (FR3).
+However, the primary focus of 55G channel modeling has been on frequencies below $6\operatorname{GHz}$ (FR1) and above $24\operatorname{GHz}$ (FR3).
 To address this gap, frequency interpolation techniques were used by 3GPP to estimate channel parameters in the $7$-$24\operatorname{GHz}$ band.
@@
 For this, true empirical characterization of the many channel parameters is required.
-3GPP channel parameters include pathloss exponents, penetration losses, shadowing effects, cross-polarization discrimination (also referred to as cross-polarization isolation), in addition to delay and angular spreads across different environments, such as indoor hotspot (InH), indoor factory (InF) and urban microcell (UMi) under various conditions, such as line of sight (LoS) and non line of sight (NLoS) with minimal outages.
-The findings of FR3 channel measurements conducted by NYU WIRELESS in [1, 4, 5] reveal that for wideband channels at $16.95\operatorname{GHz}$ in the FR3 band, the omnidirectional pathloss exponent (PLE) values, synthesized from directional channel measurements, in LoS scenarios were slightly lower than those in milli-meter wave (mmWave), indicating less signal attenuation and more of a waveguide effect over distance, e.g.
-the omnidirectional LoS PLEs in InH, UMi and InF for $6.75\operatorname{GHz}$, $16.95\operatorname{GHz}$, and $28\operatorname{GHz}$ are given in Table I, e.g.
+II-B FR3 pathloss exponents over different environments
+The findings of FR3 channel measurements conducted by NYU WIRELESS in [1, 4, 5] reveal that for wideband channels at $16.95\operatorname{GHz}$ in the FR3 band, the omnidirectional PLE values, synthesized from directional channel measurements, in line of sight (LoS) scenarios were slightly lower than those in millimeter wave (mmWave), indicating less signal attenuation and more of a waveguide effect over distance, e.g.
+the omnidirectional LoS PLEs in indoor hotspot (InH), urban microcell (UMi) and indoor factory (InF) for $6.75\operatorname{GHz}$, $16.95\operatorname{GHz}$, and $28\operatorname{GHz}$ are given in Table I, e.g.
 $1.34$ at $6.75\operatorname{GHz}$, $1.32$ at $16.95\operatorname{GHz}$, and $1.2$ at $28\operatorname{GHz}$ [1].
-In UMi, the omnidirectional LoS PLEs are $1.79$ at $6.75\operatorname{GHz}$ [1], $1.85$ at $16.95\operatorname{GHz}$ [1], and $2.02$ at $28\operatorname{GHz}$[7], which indicates slightly less loss over distance than free space channel (e.g., when PLE $=2$) and with less loss at lower frequencies.
-Although antenna patterns and gains vary widely, and higher frequencies enable the use of directional antennas that can offset channel loss in mmWave bands, 3GPP and researchers often use an omnidirectional channel model to standardize link analysis and ensure consistency in evaluations while enabling the use of any type of directional pattern to be applied to the models [9], [10].
-NLoS PLEs were found to be higher than LoS scenario as is found in all 3GPP bands as per Table I, e.g.
+In UMi, the omnidirectional LoS PLEs are $1.79$ at $6.75\operatorname{GHz}$ [1], $1.85$ at $16.95\operatorname{GHz}$ [1], and $2.02$ at $28\operatorname{GHz}$[7], which indicates slightly less loss over distance than free space channel (e.g., when PLE is $2$) and with less loss at lower frequencies.
+Although antenna patterns and gains vary widely, and higher frequencies enable usage of directional antennas that can offset channel loss in mmWave bands, 3GPP and researchers often use an omnidirectional channel model to standardize link analysis and ensure consistency in evaluations while enabling the use of any type of directional pattern to be applied to the models [8], [9].
+Omnidirectional non line of sight (NLoS) PLEs were found to be higher than LoS scenario as is found in all 3GPP bands as per Table I, e.g.
 NLoS PLE for UMi of $2.56$ at $6.75\operatorname{GHz}$ [1], $2.59$ at $16.95\operatorname{GHz}$ [1], and for InF the NLoS PLE is $1.78$ at $6.75\operatorname{GHz}$ [5], and $2.11$ at $16.95\operatorname{GHz}$ [5], the loss over distance in NLoS locations is lower (e.g., $12$ to $14.5\operatorname{dB}$ per decade less) than those observed at mmWave frequencies, e.g.
@@
 In practice, the proper comparison is to consider equal antenna aperture areas at different frequencies, where in practice the extra array gain at mmWave can overcome the additional channel loss by many dB [10].
-Referring to Table I, the PLE for LoS in UMi is $1.79$ for $6.75\operatorname{GHz}$ and $2.02$ to $2.1$ [7] for $28\operatorname{GHz}$, implying greater coverage with a $2.2\operatorname{dB}$ stronger signal per decade of distance, resulting in, $4.4\operatorname{dB}$ over $100\operatorname{m}$ and $6.6\operatorname{dB}$ stronger signal over $1\operatorname{km}$, which clearly demonstrates how a lower PLE translates into significant coverage improvement, which may be traded for higher capacity for users within a cell, a factor that will be of paramount importance to carriers when deploying $6$G technologies.
-The time delay spread of a propagation channel has impact on signaling formats, in addition to accuracy for integrated sensing and communications (ISAC) applications such as position location and multi-user synchronization methods.
-Measurements in [1], [4], [5] show that root mean square (RMS) delay spread (DS) decreases with carrier frequency, with FR3 exhibiting smaller DS values than sub-$6$ $\operatorname{GHz}$, but larger than at FR2 mmWave frequencies, e.g.
+In this work, Fig.
+2 and Fig.
+3 do not consider the reality of greater channel gain at higher frequencies, and hence overestimate the achieavable capacity of lower frequencies and underpredict higher frequency rates.
+Referring to Table I, the PLE for LoS in UMi is $1.79$ for $6.75\operatorname{GHz}$ and $2.02$ to $2.1$ [7] for $28\operatorname{GHz}$, implying greater coverage with a $2.2\operatorname{dB}$ stronger signal per decade of distance, resulting in, $4.4\operatorname{dB}$ over $100\operatorname{m}$ and $6.6\operatorname{dB}$ stronger signal over $1\operatorname{km}$, which clearly demonstrates how a lower PLE translates into significant coverage improvement, which may be traded for higher capacity for users within a cell, a factor that will be of paramount importance to carriers when deploying 66G technologies.
+II-C FR3 delay spreads, angular spreads & penetration losses
+The DS of a propagation channel has impact on signaling formats, in addition to accuracy for integrated sensing and communications (ISAC) applications such as position location and multi-user synchronization methods.
+Measurements in [1], [4], [5] show that root mean square (RMS) DS decreases with carrier frequency, with FR3 exhibiting smaller DS values than sub-$6$ $\operatorname{GHz}$, but larger than at FR2 mmWave frequencies, e.g.
 $14\operatorname{ns}$ and $12.7\operatorname{ns}$ LoS omnidirectional RMS DS at $6.75\operatorname{GHz}$ and $16.95\operatorname{GHz}$, respectively for InF.
 The range of RMS DS was found to be $12.7$-$37.7\operatorname{ns}$ across InF, UMi and InH.
-DS values observed in the field indicates that multipath components are closer in time at FR3, suggesting limited temporal dispersion which could favor high-speed, low-latency communications and more accurate position location and timing methods in dense environments.
+DS values observed in the field indicates that multipath components are closer in time at FR3, suggesting limited temporal dispersion which could favor high-speed, low-latency communications and more accurate location and timing methods in dense environments.
 Furthermore, the measurements in [1] found that higher FR3 frequencies, i.e.
-$16.95\operatorname{GHz}$, exhibited a narrower RMS angular spread compared to the higher FR1 frequency at $6.75\operatorname{GHz}$, hence indicating fewer, more focused multipath components, which is beneficial for spatial multiplexing as it reduces interference between signal paths and allows more precise beamforming.
+$16.95\operatorname{GHz}$, exhibited a narrower RMS AS compared to the higher FR1 frequency at $6.75\operatorname{GHz}$, hence indicating fewer, more focused multipath components, which is beneficial for spatial multiplexing as it reduces interference between signal paths and allows more precise beamforming.
 Regarding material penetration, losses were consistently higher at $16.95\operatorname{GHz}$ in FR3 than at lower frequencies, with losses dependent on material type and polarization configuration, confirming the inherent limitations of FR3 in penetrating certain materials, such as low-emissivity (IRR) glass and concrete [1], yet able to allow more penetration that at FR3 mmWave [1].
-When considering the design and comparison of air interface standards, such as the 3GPP TR 38.901 statistical channel model, the channel measurements [4, 5] conducted at new FR3 frequencies suggest that no major changes are needed in terms of numerology compared to existing 5G- new radio (NR) protocols; however, the beam management may need modifications due to the different temporal and spatial statistics and the different number of antenna elements at both gNodeB (gNB) and user equipment (UE).
-In particular, directional links rely on precise angular alignment of transmitter and receiver beams, a process facilitated through beam management operations.
-For next-generation cellular networks, mechanisms are being discussed in 3GPP technical documents that address beam management, such as R1111-$2409673$ in November $18-22$ as part of 3GPP technical specification group (TSG) radio access network (RAN) working group (WG1), in addition to other 3GPP contributions such as R1111-$2409710$, R1111-$2409741$ and more.
-Directional combining in beam management dramatically enhance link gains and coverage distances, as first shown in [12], thereby maintaining acceptable communication quality when beam combining is available – something that both FR2 and FR3 will support.
-However, as shown in [4] , as the carrier frequency increases, the angular spread tends to decrease.
-In addition, the empirical RMS angular spread of arrival (ASA) at both $6.75\operatorname{GHz}$ and $16.95\operatorname{GHz}$ was noticeably lower than the 3GPP model predictions [4].
-Factors such as mobility, device rotation, and wide angular spread channels can impact beam alignment.
-In dynamic scenarios, beam measurements can quickly become outdated, leading to inaccurate beam selection and performance degradation, which further impacts signal-to-interference-plus-noise ratio (SINR) and spectral efficiency (SE).
-The newly reported angular spreads in [4] implies that modifications are required for beam management in the new goldilocks FR3 mid-band spectrum.
-For sensing applications in ISAC, this also means that precise and accurate beam alignment is needed for target detection, localization and tracking.
-In addition to channel modeling, the call for new metrics and figures-of-merit, such as the Waste Factor, offer potential for energy efficiency optimization in $6$G wireless systems.
-II-B FR3 frequency-dependent features and losses
-As discussed in the previous section, empirical field data [5] in InF (dense-scatterer environment) indicate frequency-dependent trends for DS and angular spread (AS).
-This is not suprising, as frequency-dependence propagation features were measured and revealed in [13] for both outdoors and indoors empirically discovered in what eventually became the FR3 bands of $28$ and $73\operatorname{GHz}$.
+II-D FR3 frequency-dependent features and losses
+Frequency-dependent propagation features were measured and revealed in [11] for both outdoors and indoors empirically discovered in what eventually became the FR3 bands of $28$ and $73\operatorname{GHz}$.
 For modeling the channel path loss over distance, extensive empirical measurements show the close-in free space model with a frequency-weighted path loss exponent (CIF) is more suited for indoors [11], and extends the CI model which has a physics-based close-in free space reference distance as a leverage point for the slope of the exponential path loss, while incorporating the frequency dependence feature of path loss observed indoors.
-The multipath temporal and spatial statistics at the lower part of the FR3, e.g.
-$7\operatorname{GHz}$, varies considerably compared to the higher part of FR3, e.g.
-$16$-$24\operatorname{GHz}$, which confirms less multipath dispersion in both time and space at higher mmWave frequencies [1].
 Perhaps most importantly, rain attenuation and foliage losses vary significantly across the FR3.
-Following ITU-R recommendation (P.838838838838-$3$), the specific attenuation at $7\operatorname{GHz}$ for heavy rain (corresponds to rain rate of $8\operatorname{mm/hr}$) is $0.04\operatorname{dB/km}$, whereas at $24\operatorname{GHz}$, the specific attenuation is $1.16\operatorname{dB/km}$, i.e.
+Following ITU-R recommendation (P.838838-33), the specific attenuation at $7\operatorname{GHz}$ for heavy rain (corresponds to rain rate of $8\operatorname{mm/hr}$) is $0.04\operatorname{dB/km}$, whereas at $24\operatorname{GHz}$, the specific attenuation is $1.16\operatorname{dB/km}$, i.e.
 a gap of $1.16\operatorname{dB}$ difference over $1\operatorname{km}$.
@@
 However, the link can be improved to extend the transmitter’s coverage area via coherent multi-beam combining [12].
-III Spectrum Agility & Coverage-Rate Tradeoffs
-III-A Spectrum Agility for Coverage-Rate Tradeoffs
-Under a constant EIRP with shrinking aperture assumption, the lower FR3 frequencies provide better coverage, as detailed in Section II.
-If instead the antenna physical area is held fixed, the extra array gain available at mmWave more than compensates for the additional free space pathloss.
-A fundamental question arises: What mechanism can optimally balance rate-coverage tradeoffs ?
-A particularly interesting ray tracing simulation was conducted in [14] to study the coverage and rate statistical behaviour across the FR3 spectrum.
-In terms of rate, with no blockages, and as expected, it was shown that $24\operatorname{GHz}$ usage dominates in terms of rate for outdoors.
-Interestingly, when a 3GPP blockage model B was used, it was shown that there was a fallback need towards $12\operatorname{GHz}$ and $6\operatorname{GHz}$ for maximum rate.
-But even more, and making use of the the comprehensive and detailed penetration models and findings from NYU WIRELESS [1], it is observed that while materials like glass and wood exhibit relatively permeable characteristics at FR3 frequencies, others, e.g.
-low-emissivity glass, are not permeable.
-A particularly intriguing observation is that concrete, which is relatively permeable at sub-$6$ $\operatorname{GHz}$ frequencies, becomes completely impermeable at FR3.
-Additionally, as reported in [14], the indoor user rate distribution indicates that users operating at $24\operatorname{GHz}$ will need to fallback to lower frequencies, which means that, if we really want to leverage the large FR3 spectrum, frequency hopping mechanisms between FR3 bands should be designed to account for the various blockages and penetration losses.
-The diversity and hopping mechanism naturally pushes forward the need for a spectrum agility in the FR3.
-Besides, the agility also inherently secures certain transmissions from malicious users and eavesdroppers.
-A natural question then to ask would be what hopping mechanism should be adopted between the FR3 bands in order to be spectrally efficient and agile?
-III-B Incumbents: Coexistence & Opportunity
-The primary challenge of utilizing FR3 for $6$G lies in regulatory issues.
-The spectrum is densely populated with both civilian and government incumbents for various applications beyond fixed and mobile wireless access, like meteorology, radio astronomy, and maritime radio navigation.
-In contrast to $5$G, one would need to pay special care towards such incumbents.
-Even if regulators reach consensus on spectrum availability and licensing frameworks, the most technically demanding aspect will be determining how to share this spectrum without interfering with existing users.
-Note that satellite users are already spectrum-sharing amongst themselves.
-Vice-versa, we can also consider a fundamental question that is: are commercial services and malicious users able to harm satellites?
-A possible approach would be through interference nulling to guarantee protection of satellite uplink (UL) communications without sacrificing much the terrestrial downlink (DL), which translates into a greater chance of co-existence.
-Sharing spectrum efficiently while avoiding interference, especially in cases where users operate close to one another, is a technical challenge that requires effective coordination mechanisms.
-Relocating incumbent users can be expensive, often eclipsing the auction value of the spectrum itself.
-Inspired by spectrum refarming mechanisms, dynamic spectrum sharing (DSS) can ensure backward compatibility with older generations in a dynamic way.
-More specifically, DSS allows service providers to bypass a static cut between different generations in frequency, allowing for wiser spectral utilization.
-Similarly, DSS can allow a dynamic smart way of sharing spectrum between $6$G users and incumbents within the band, but this would be very challenging.
-IV Massive MIMO, Compact Arrays & Power Considerations
-Doubling the frequency implies a loss by a factor of $4$ in Watts.
-However, this loss can be regained by filling the same effective area, covering the space, which also means more compact (but not bigger) arrays, which is referred to as antenna gain.
-Given half a wavelength spacing, then multiplying the carrier frequency by $N$ necessitates $N^{2}$ more antennas to cover the aperture area.
-For FR3, transitioning from $3.5\operatorname{GHz}$ to $7\operatorname{GHz}$ faces a $6\operatorname{dB}$ free space path-loss, but also a $6\operatorname{dB}$ antenna gain assuming the same aperture size, enabling similar coverage to lower frequencies with higher efficiency.
-Meanwhile, the coherence distance shrinks with higher frequencies, which implies that the minimum distance over which signals are uncorrelated in phase and amplitude also decreases, giving rise to the need of packing up more antennas in a given aperture.
-The challenge remains in the NLoS case, especially for neutral host, where there are much more severe penetration losses at higher frequencies compared to lower counterparts, as mentioned in Sections II and III, in addition to diffraction losses which scale as $10\log f$.
+It is critical to realize that the 3GPP channel modeling efforts in TR 38.901 [6] have up until now ignored the frequency-dependent channel characteristics for most temporal/spatial statistics between 66-$100$ GHz, hence these real-world data are vital for informing a new version of the 3GPP channel model in coming years [4, 5].
+Figure 2: Multi-frequency ISAC trade-offs between low and high frequencies.
+Figure 3: Rate-coverage analysis over 77 GHz, $14$ GHz, $18$ GHz and $24$ GHz.
+TABLE II: Pros and cons of exploiting the 77-$24$ GHz FR3 spectrum for 66G
+Pros / Opportunities Cons / Challenges & Research gaps
+Wider bandwidth than FR1 but lower pathloss than FR2.
+Balances coverage with capacity: outdoor-to-indoor reach that mmWave lacks, but higher throughput than FR1.
+The centimeter wavelengths enables large antenna, unlocking extreme massive MIMO (mMIMO).
+Narrower beams than FR1 and richer multipath than FR2 make FR3 attractive for ISAC.
+Compact high-gain arrays possible on drones.
+Compact arrays with mutual coupling for tightly coupled colinear antenna arrays can expand bandwidth [13].
+Rain/atmospheric loss lower than mmWave.
+”Bridge band” for sensing, which helps in mitigating Doppler mismatch between FR1+FR3 or FR3+FR2.
+High-resolution ISAC: 100100400400 MHz chunks, good for multi-target tracking and imaging.
+Multi-band capabilities yielding a diverse view of the environment.
+Coverage–capacity sweet-spot: 10 GHz link budget still supports O2I service without mmWave densification.
+Mid-band inheritance: more loss than FR1, less available bandwidth than FR2.
+phase noise (PN) at higher frequencies leads to signal-to-noise ratio (SNR) degradations.
+Aggregating separated sub-bands increases ADC and algorithmic complexity; frequency-dependent channels need characterization.
+RF front-end parts (including power amplifiers, analog-to-digital converters, and mixers) covering 772424 GHz are immature.
+Antenna design for entire FR3 is a challenge.
+Unified channel models for terrestrial, NTN and ISAC links still ongoing.
+Spectrum coexistence and aggregation policies demand new multidimensional management schemes.
+Crowded spectrum: lower FR3 hosts fixed links & DoD users, in addition to X-band for synthetic aperture radar (SAR).
+Upper FR3 overlaps Ku-band.
+Sharing needed.
+Incumbent protection needs accurate sensing of satellite terminals.
+Open research gaps: unified near/far-field models, extremely large aperture array (ELAA)-aware ISAC theory, learning-based spectrum sharing.
+III ITU enhancements via FR3
+Peak data rates & bandwidth requirements
+Shannon’s formula of capacity tells us that doubling the bandwidth offers higher rates than doubling the power.
+To keep up with the momentum of different cellular generations, carrier bandwidth should be extended.
+Indeed, code-division multiple access (CDMA) in 33G used a spectrum block (SB) of $5\operatorname{MHz}$, whereas long-Term Evolution (LTE) uses $20\operatorname{MHz}$.
+For 55G FR1, the SB is $100\operatorname{MHz}$.
+Moving forward, 66G must open SBs that are at least $400\operatorname{MHz}$ wide to meet the increasing capacity demands of applications like augmented reality (AR)/ virtual reality (VR).
+For that, if $3$ carriers in FR3 are supported, one can go up to $1.2\operatorname{GHz}$ of possible non-contiguous bandwidth.
+This numerology seems very reasonable to keep up with the ITU 66G vision to achieve peak data rates of $50\operatorname{Gbps}$ to $200\operatorname{Gbps}$ depending on the scenario.
+A $1024$- quadrature amplitude modulation (QAM) constellation offers $10\operatorname{bits/DoF}$.
+With $12$ antennas providing $12$ spatial degrees-of-freedom (DoF) and $1.2\ \operatorname{GHz}$ overall bandwidth, a peak data rate of $144\ \operatorname{Gbps}$ is achieved.
+To meet and exceed the IMT-2030 requirement of $200\ \operatorname{Gbps}$, the system can:
+  • Increase spatial streams (e.g., $16$ antennas to offer $192\operatorname{Gbps}$)
+  • Aggregate more carrier bandwidth additional $400\operatorname{MHz}$ to achieve $192\operatorname{Gbps}$.
+  • Combine the above to reach $256\operatorname{Gbps}$, surpassing the IMT-2030 vision of $\boldsymbol{200\operatorname{Gbps}}$.
+Location accuracy
+The multiband results in Fig.
+2 show that scaling carrier frequency and bandwidth by $4f_{\mathrm{ref}}$ (with $f_{\mathrm{ref}}=8\ \operatorname{GHz}$) yields a Cramér-Rao bound (CRB) for delay estimation below $10^{-3}$ at $17$ dB SNR, corresponding to sub $\boldsymbol{10}$ cm location accuracy, fully meeting the IMT-2030 target.
+When penetration loss at high bands is an issue, hopping to lower frequencies can attain better spectral efficiency (SE).
+Coverage-rate tradeoffs & Spectrum agility
+In Fig.
+3, we run a coverage-rate analysis for four distinct frequencies at FR3.
+We have allocated more bandwidth to higher frequencies, i.e.
+$7$ GHz, $14$ GHz,$18$ GHz and $24$ GHz are allocated $100$ MHz, $200$ MHz, $300$ MHz, and $400$ MHz, respectively.
+Analysis shows that up to $200\operatorname{Mbps}$, lower frequencies dominate, with $7\operatorname{GHz}$ sustaining $200\ \operatorname{Mbps}$ for $30\%$ of the time.
+Beyond $200\ \operatorname{Mbps}$, the optimal band shifts to $18\ \operatorname{GHz}$ (up to $575\operatorname{Mbps}$), and again at $850\operatorname{Mbps}$, demonstrating the need for spectrum agility in FR3.
+Dynamically hopping between FR3 frequencies can maintain optimal rate-coverage tradeoffs and deliver IMT-2030 user experienced data rate requirements for $\mathbf{\geq 1\ Gbps}$ peak and high average rates.
 […]
-It is worth noting that new approaches have emerged for the design of ultra-wide band (UWB) antenna arrays whereby mutual coupling (MC) is intentionally introduced between different elements to expand the bandwidth.
-As far as massive MIMO (mMIMO) is concerned, it was shown that compact mMIMO can open extra degrees-of-freedom in frequency domain [15].
-Specifically, MC between array elements can be leveraged in tightly coupled colinear antenna arrays to expand the bandwidth.
-While interesting to attain higher rates, how can sensing be performed under intentionally induced MC in compact mMIMO designs, especially given its frequency dependence across the UWB?
-IV-B Indoor reflection modeling for extended targets
-The reflections arising from a single target share a common relation simply because they arise from the same target, i.e.
-extended target (ET).
-ETs can reveal information besides location, e.g.
-size, shape, and orientation, which can enable use cases like gesture recognition.
-In fact, ET modeling is particularly relevant to indoors owing to its property of rich multipath reflections.
-Therefore, more modeling efforts should be conducted to characterize ET on lower and higher counterparts of FR3.
-Even more, as alluded to in Section IV-A, and besides UWB, mMIMO can allow resolving ET paths in spatial domain.
-IV-C mMIMO architectures: Fully-digital vs. hybrid
-As number of antennas grow large to support various FR3 ISAC applications, which will heavily depend on SE and multi-band operations, a fully-digital scheme may seem complex, due to high power consumption issues.
-For ISAC, one can consider a fully digital array employing low-resolution digital-to-analog and analog-to-digital converters to trade-off quantization noise with power consumption.
-Therefore, we can ask what is an optimal fully digital design that can balance sensing tasks with reasonable power consumption ?
-In contrast, hybrid beamforming seems like a good tradeoff to alleviate the power consumption problem, by trading-off some beamforming and beam selection gains, which, in turn has negative reverberation effects on the received SINR and SE for communications, in addition to detection and localization accuracy for sensing.
-Therefore, what is the optimal analog beamforming number of sub-panels to strike such trade-offs in hybrid design for sensing and communication (S&C)?
-IV-D Coherence time across FR3
-Higher frequencies naturally reduce the coherence time, which in turn increases the complexity of channel estimation and makes spatial multiplexing more challenging.
-An inherent property of FR3 is that the coherence time can significantly vary across FR3, especially in mobile and highly dynamic environments.
-The coherence block shrinks at higher frequencies, hence the channel outdates faster.
-In turn, this can limit the duration available for multiplexing data, which can also limit the number of terminals that can be multiplexed simultaneously due to constraints on orthogonal pilot sequences fitting within the coherence time.
-Within the coherence time, we also know that the achievable rate grows as $\min\left(N_{T},N_{R}\right)\log_{2}(1+\operatorname{SNR})$ per unit of frequency.
-Hence, mMIMO can allow for more data multiplexing in an environment where less coherence time is available, especially at the higher end of FR3 bands.
-However, special attention is needed for channel state information acquisition.
-A question here could be: Should higher FR3 frequencies be dedicated to less mobile environments, while lower FR3 bands to highly dynamic environments ?
-IV-E RF frontend efficiency and array integration limits
-Energy efficiency of power amplifiers can becoem a challenge in the FR3 band, so transmit side losses can limit the link budget advantages.
-Particularly, packing large numbers of elements into a fixed aperture soon hits practical barriers due to thermal dissipation, EIRP limits, and fronthaul capacity.
-Therefore, from array design and RF frontend perspective, how can future FR3 radios achieve high energy efficiency in their power amplifiers while packing dense antenna arrays under real world thermal, EIRP, and processing limits, without giving up the very coverage-rate gains the band is meant to provide?
-Waste factor can serve as a fundamental metric towards the design.
-V Multiband sensing for FR-3 ISAC
-ISAC use cases planned for FR3 mean that channel models must support S&C functions.
-Multi-band sensing for future systems can be realized through non-contiguous aggregation, whereby different frequency blocks are spread across different bands or parts of the FR3 spectrum.
-Next, we provide an insightful analysis of some aspects of multiband sensing.
-As frequency increases, there is typically more spectrum available, and more antennas can be packed per unit of space.
-Thus, we assume that the bandwidth and antenna gain is proportional to the center frequency.
-It can be verified that when the frequency is doubled, the signal-to-noise ratio (SNR) and the Cramér-Rao bound (CRB) drop by a factor of $2$.
-An example is provided in Fig.
-2.
-The drop of SNR at higher frequencies shifts the waterfall region, and the lower frequencies have more margin to accurately start estimating the delay as predicted by the CRB.
-Conversely, since the CRB of the higher frequency system is smaller, it means that after the waterfall region, the higher frequency achieves a more accurate estimate.
-Figure 2: Multi-frequency ISAC trade-offs between low and high frequencies.
-Our analysis reveals two insights for leveraging multi-band sensing.
-The first is to obtain estimates individually and combine via a maximum ratio combining type approach.
-Specifically, one can assign more weights to estimates obtained at higher frequencies.
-The second is relevant for bistatic sensing, where we can rely on higher FR3 frequencies to obtain synchronization via LoS path due to the need of accurate timing, but then we can tolerate to fallback to lower frequencies, which we know can cover more targets.
-Based on this scheme, a good time difference of arrival location-based method can be developed leveraging the entire multi-band information on FR3.
-VI Key FR3 features in $6$G Standards
-This section explores key modifications envisioned for next-generation $6$G cellular standardization to unlock the full potential of spectrally agile FR3 radios.
-Specifically, we examine both 3GPP innovations—focusing on radio frame enhancements tailored for spectral agility—and new O-RAN-based features, particularly in O-RAN open radio unit (O-RU) designs, that facilitate intelligent spectrum utilization.
-3GPP
-The challenges posed by FR3, e.g.
-wider bandwidths, non-contiguous carriers, increased propagation losses, and dynamic spectrum environments, necessitate innovations in frame design to ensure low-latency, high-reliability, and efficient multi-band operations.
-Below, we outline key advancements required to enhance spectral agility, synchronization, and resource management in next-generation wireless systems:
-  • Multi-band aggregation, i.e. combining separate, non-contiguous spectrum blocks to optimize the balance between performance and complexity, enhancing both S&C capabilities.
-  • Dynamic slot configuration for multi-band coordination, which implements AI-driven adaptive slot structures that dynamically adjust frame configurations across multiple bands.
-The dynamic strategy ensures optimized latency, throughput, and reliability by tailoring subframe allocations to each band’s propagation characteristics.
-  • Multi-band pilot design, which aims at designing hybrid pilot structures that combine common and band-specific pilots while employing sparse allocation, frequency sharing, and adaptive density.
-Multi-Band pilot design can be realized via AI-driven optimization and compressed sensing techniques to minimize redundancy and intelligently manage pilots for greater efficiency in multi-band environments.
-  • Agile timing advance, which aims at developing new timing advance strategies to ensure seamless switching and synchronization across carriers in different frequency bands, which is critical for achieving low-latency and high-reliability performance in $6$G.
-  • Multi-carrier synchronization and scalable numerology, meaning that it is vital to introduce enhanced scalable numerology to support multi-carrier synchronization across diverse FR3 bandwidths, which further allows adaptive subcarrier spacing and multi-band coordination without excessive signaling overhead.
-[…]
-From RAN perspective, the open radio access network (O-RAN) RAN intelligent controller (RIC) plays a crucial role in enabling spectral agility and intelligence by leveraging real-time (RT) and non-RT network optimizations.
-Through the use of xApps (near-RT RIC applications) and rApps (Non-RT RIC applications), the RIC can dynamically manage spectrum resources, adapt to changing conditions, and optimize performance in FR3.
-Hereafter, we examine how spectral agility and spectral intelligence can be achieved through the RIC.
-Spectral Agility: The ability of a network to dynamically reconfigure spectrum usage based on real-time conditions, interference levels, traffic demands, and regulatory constraints can be achieved through the following mechanisms:
-  • Dynamic carrier aggregation & multi-band adaptation: The Near-RT RIC can monitor real-time spectrum conditions and trigger carrier aggregation adjustments across multiple FR3 bands, which enables on-the-fly switching between licensed, shared, and unlicensed spectrum resources, optimizing throughput and reducing congestion.
-  • Intelligent spectrum sensing and sharing: By integrating AI/ML models within rApps, the Non-RT RIC can analyze historical spectrum usage patterns and proactively allocate spectrum to different cells or operators, thus enabling dynamic spectrum access and coexistence with satellite, radar, or other wireless services operating in FR3.
-Spectral Intelligence: Efficiency, reliability, and adaptability of spectrum can be achieved through intelligent decision-making and automation.
-  • Beamforming, channel estimation, and interference management: xApps in the Near-RT RIC continuously monitor spectrum conditions and adjust beamforming strategies for multi-user MIMO in FR3.
-Advanced channel estimation and interference prediction can help minimize co-channel and adjacent-band interference, ensuring optimal signal quality.
-  • Multi-band O-RU control for distributed MIMO: The Near-RT RIC can coordinate multi-band O-RUs across FR3 to enable distributed MIMO architectures.
-Leveraging O-RAN’s open fronthaul interface, it ensures tight synchronization across spatially separated RUs, maximizing spectral efficiency.
-  • Power efficiency: Other key innovations that can be implemented by leveraging the RIC are adaptive sampling and dynamic range control.
-By adjusting these parameters based on real-time spectral conditions, FR3 O-RUs reduce unnecessary power consumption and enhance operational efficiency.
-VII Conclusions
-This paper outlined a vision for $6$G in the FR3 spectrum, exploring its upper mid-band potential, key challenges, and necessary standardization changes.
+Doubling the carrier frequency, e.g.
+$3.5$ to $7$ GHz, introduces a $6$ dB free space pathloss and shorter coherence distances, requiring $N^{2}$ more half-wavelength elements to fill the same aperture, but this is counterbalanced by an equal $6$ dB antenna gain, while higher frequency NLoS links still face larger penetration and diffraction losses.
+mMIMO can aid in achieving the spectrum efficiency specified in IMT-2030, given a preliminary study on the number of antennas required.
+For IMT-2030 connection density, besides resorting to lower bands for coverage, uplink (UL) densification, where operators can add more transmission and reception points, already supported in 3GPP Rel-16, which are physical base stations within a wireless system that act as nodes for transmitting and receiving signals to improve UL coverage and capacity by attaining spatial diversity at the macro level so that if one path is blocked, an alternative can be used instead.
+TABLE III: Comparison among FR1, FR2, and FR3
+FR Deployment feasibility Coverage Rate Hardware Implications PLE DS AS
+FR1 Outdoors Wide Low Low cost Low High Broad
+FR2 Indoors & backhaul Narrow High High complexity High Low Narrow
+FR3 Outdoors & indoors Medium Mid Flexible Moderate Moderate Moderate
+IV Challenges for commercial deployment
+In this section, we highlight some challenges for commercial deployment supporting FR3, which include
+Non-contiguous bandwidth & dynamic spectrum environment
+Non-contiguous carrier aggregation requires higher resolution ADCs and digital-to-analog converters and more sophisticated baseband algorithms, thereby increasing processing complexity and power consumption.
+Enabling dynamic spectrum access and coexistence with satellite, radar, or other wireless services operating in FR3 is crucial.
+Phase noise (PN) at higher frequencies
+PN arises as noise occurring from random and rapid phase fluctuations of a given signal.
+Hence, PN in both clocks and oscillators can create impact the performance, especially at higher frequencies.
+In sampling, for a given jitter error, higher frequency signals will suffer from higher phase errors, therefore degrading the clock SNR.
+In analog, a given amount of time jitter also leads to higher PN at higher frequencies.
+Over FR3, an SNR degradation caused by PN translates to a loss of $10.7\operatorname{dB}$.
+Spectrum coexistence
+Lower FR3 bands overlap fixed-link and defense allocations, while upper FR3 intersects Ku-satellite services, necessitating accurate incumbent sensing and multidimensional spectrum-sharing policies.
+RF frontend, impairments & linearity
+PAs, mixers, and tunable filters capable of broadband operation over FR3 are not yet available at commercial scale.
+In addition, RF front ends, have to meet stringent adjacent channel leakage power ratios over multiple sub-bands within size constraints of user equipments.
+Moreover, carrier frequency offset and inphase/quadrature imbalance vary across bands, requiring new estimation and compensation techniques.
+Furthermore, easily tunable filters to operate smoothly across different FR3 bands is a crucial aspect for multi-band operations in order to leverage the full capability offered by FR3.
+Also, operating in FR3 necessitates advanced electronic components, such as ADCs and PAs, which must support minimal distortion and maintain linear performance over wider bandwidths.
+Additional challenges can be found in Table II.
+V Open Questions & Potential Solutions
+Question 1: How do recent channel-measurement findings at new FR3 mid-band frequencies affect strategies in 5G-NR (and beyond), and what protocol or metric modifications are being considered within 3GPP to maintain link quality and energy efficiency (EE)?
+The newly reported ASs in [4] implies that modifications are required for beam management in the new goldilocks FR3 mid-band spectrum.
+For sensing applications in ISAC, this also means that precise and accurate beam alignment is needed for target detection, localization and tracking.
+In addition to channel modeling, the call for new metrics and figures-of-merit, such as the Waste Factor, offer potential for EE optimization in 66G wireless systems.
+Question 2: What is an optimal fully digital design that can balance sensing and communication tasks with reasonable power consumption ?
+Hybrid beamforming seems like a good tradeoff to alleviate the power consumption problem, by trading-off some beamforming and beam selection gains, which, in turn has negative reverberation effects on the received signal-to-interference-plus-noise ratio (SINR) and SE for communications, in addition to detection and localization accuracy for sensing.
+Question 3: How to deal with the difference in coherence time between the lower and upper ends of FR3 ?
+The lower end frequencies of FR3 can be allocated to more highly dynamic environments, whereas the higher end to less mobile environments.
+Question 4: How to promote SE and agility?
+Frequency hopping is interesting in NLoS with severe blockages emanating from building materials, where gains depend on the hopping strategy.
+However, to study hopping gains, practical antenna gains based on aperture size that fits a BS and UE must be used for assessment, as higher band systems employ higher antenna gains than lower bands.
+Typical additional link budget gains are about $26$ dB at the BS and $13$ dB at the UE for $28$ GHz, compared to $10$ dB and $6$ dB respectively at $7$ GHz.
+Using realistic gains, rather than the omni-directional assumptions in TR 38.901, will yield much more reasonable performance curves, often showing only marginal improvements for mid-band compared to high-band in such scenarios.
+Question 5: Are commercial services and malicious users able to harm satellites on FR3?
+Yes.
+Terrestrial networks can unintentionally disrupt satellites via co-channel interference or even collision risk from dense constellations.
+Even more, malicious users go further, deliberately jamming, spoofing, hacking or physically attacking spacecraft, potentially denying service leading to seizing control.
+For that, mitigation should rely on techniques such as interference nulling and dynamic spectrum sharing (DSS), yet none offers total immunity.
+Consequently, robust coordination rules and continuous security monitoring are essential to ensure satellite safety.
+Also, harmonious operation with incumbents can be attained via dynamic spectrum access with real-time interference nulling beams and incumbent sensing capabilities.
+Question 6: How does FR3 propagation characteristics, effective bandwidth, and antenna array size in the FR3 band bound the positioning error of received signal strength (RSS)-, angle-, and time-based cellular localization with respect to the IMT-2030 111010 cm target?
+RSS
+-based systems alone, even with several synchronized BSs may not be enough to satisfy the stringent 111010 cm target set by IMT under typical high pathloss and shadowing.
+Angle-based positioning with large uniform arrays may achieve the required accuracy in LoS but fails in dense urban multipath, revealing strong sensitivity to AS in this spectrum.
+Therefore, joint multiband processing for time-based and angle-based positioning algorithms may be needed to attain the 111010 cm target.
+Part of future work should be able to assess the exact localization accuracy over FR3 under its diverse propagation characteristics.
+Question 7: What new propagation traits, never captured in FR1/FR2, can FR3 reveal?
+One fundamental feature of operating in the FR3, particularly relevant for sensing, is the phenomenon of diffuse scattering [14], leading to dense multipath components (DMC).
+Surfaces and objects are seen differently over FR3 depending on the wavelength of the incident wave [14].
+Within FR3, and for ISAC applications, a target/object interacting with a channel at the lower end of FR3, will naturally have different diffuse scattering properties than that at the upper end of FR3, even when considering the second-order moments of that channel.
+In turn, the DMC statistics can offer a distinctive feature of identifying and classifying different targets across FR3.
+In addition, FR3 will accommodate hybrid near-field and far-field beamforming and algorithms as the Fraunhofer distance varies significantly over the lower/upper FR3 spectrum.
+Question 8: What key 3GPP and O-RAN innovations enable spectral agile & spectral intelligence FR3 operation in 6G networks?
+  • Multi-band aggregation, which stitches together non-contiguous FR3 blocks so radios can treat scattered spectrum as one wide virtual channel, boosting both sensing and data throughput.
+  • Dynamic slot configuration using AI, which can reshape frame structures on the fly across multiple bands to hit the sweet spot between latency, reliability, and capacity.
+  • Multi-band pilot design, which compresses and shares pilots intelligently via sparse allocation, in addition to frequency sharing and adaptive density.
+Multi-Band pilot design can be realized via AI-driven optimization and compressed sensing techniques to minimize redundancy and manage pilots for better efficiency in multi-band environments.
+  • Agile timing advance and scalable numerology, which keep devices tightly synchronized as they hop among carriers with different sub-carrier spacings, sustaining low-latency links.
+In addition, an AI-guided RAN intelligent controller (RIC) orchestrates carrier aggregation, beamforming, interference mitigation, and power-adaptive O-RUs so the full FR3 spectrum is exploited efficiently and energy-consciously.
+Near-RT xApps run inside the RIC’s real-time loop, reacting within milliseconds to spectrum-usage shifts to retune carriers and beams, giving the network reflex-like agility.
+Regarding legacy RAN, even where incumbent non-O-RAN platforms dominate, 66G spectrum agility can still be achieved.
+Operators can for example: $(i)$ retrofit wideband and multiband remote radio units and software-defined basebands already supported by most LTE and 55G sites; $(ii)$ activate standards-based features such as DSS, expanded carrier aggregation, in addition to integration with other important technologies such as self-organizing networks and auto refarming.
+$(iii)$ In addition, an AI-driven, closed-loop assurance layer that ingests real time RAN data, processes it through a machine-learning pipeline, and then issues automated network actions allows carriers to be split, resized, or retuned on the fly whenever slice traffic or service requirements change.
+Such approaches can deliver O-RAN like features without a disruptive replacement of legacy RAN hardware.
+VI Conclusions
+This paper outlined a vision for 66G in the FR3 spectrum, exploring its upper mid-band potential, key challenges, and necessary standardization changes.
 We highlighted FR3’s advantages in capacity, coverage, and SE, while emphasizing the impact of its channel characteristics on beam management.
 Based on our FR3 measurements, we underscored the need for frequency hopping to mitigate blockages, maximize SE, and incumbent coexistence, reinforcing the importance of spectrum agility.
-We examined the potential of mMIMO for FR3 and discussed design challenges, particularly in the uplink.
+We examined the potential of mMIMO for FR3 and discussed design challenges, particularly in UL.
 For sensing, we analyzed performance bounds for FR3 ISAC multiband sensing.
-Additionally, we detailed key $6$G standardization features, including 3GPP radio frame modifications and O-RAN-based innovations, using RIC-driven xApps and rApps for intelligent spectrum management.
+Additionally, we detailed key 66G standardization features, including 3GPP radio frame modifications and O-RAN-based innovations, using RIC-driven xApps and rApps for intelligent spectrum management.
 However, open questions remain regarding FR3 propagation characteristics, PHY design, and regulatory considerations.